OpenStudy (anonymous):

Express answer in exact form. A regular hexagon with sides of 3" is inscribed in a circle. Find the area of a segment formed by a side of the hexagon and the circle. (Hint: remember Corollary 1--the area of an equilateral triangle is 1/4 s2 √3.)

OpenStudy (anonymous):

To find the area of a segment, we subtract the area of a triangle from the area of a sector. I understand this, but when i try this i get a certain answer, and the answer i am supposed to give should be written as: \[A=(__\pi-_√_)inches^2\]

OpenStudy (anonymous):

Can you draw it.....

OpenStudy (anonymous):

Draw what? I am not given a shape or illustration.

OpenStudy (anonymous):



OpenStudy (anonymous):

The area of the segment = \[\frac{ 1 }{ 6 } * \pi * 3^{2} - \frac{ \sqrt{3} }{ 4 } * 3 ^{2} \]